OpenStudy (anonymous):

The formula r=2sprt(5L) can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of the length L, in feet. How far will a car skid at 30 mph?

OpenStudy (anonymous):

r=2sqrt(5L) 50=2sqrt(5L) 50/2=25=sqrt5L (25)^2=5L [(25)^2]/5=L L=125 feet